# Simulating quantum computation: how many “bits” for “it”?

### v2

@@ I Introduction
This description of quantum computation with magic states grew out of the analysis of Wigner function negativity as a precondition for a quantum computational speedup, a research programme that started with Refs. [3] and [4]. Specifically, in Ref. [4] it was established that negativity in Gross’ Wigner function [5, 6] is required for a quantum speedup, under the condition that the Hilbert space dimension is odd. Analogous results were subsequently established in even dimension, specifically for rebits [7] and qubits [8]. However, in the end it turned out that once sufficiently general (quasi)probability functions are admitted, there is no need for any negativity at all [2, 9]. Universal quantum computation in the magic state model can be described by repeated sampling from a generalized phase space whose points are labeled by the vertices of the $\Lambda$ polytopes. This process essentially resembles a random walk, with the complication that the transition function changes from one time step to the next and can depend on the prior sampling history.
Figure 1: Quantum computation with magic states (a), and its simulation based on $\Lambda$ polytopes (b). (a) QCM consists of preparing a quantum register in a magic state $|M\rangle^{\otimes n}$, followed by a sequence of Pauli measurements. This requires a device Prep to deliver the magic states to the quantum register QR, and a classical register CR to store the previous measurement record s, a classical side computation to identify the label $a(t)$ of the Pauli observable measured in step $t$, and a measurement device M to perform the measurements and to output the corresponding results $s(t)$. (b) The overall structure of the classical simulation is the same, but with the components modified. Prep is replaced by a first sampler $\mathsf{\Sigma}$ that samples from the phase space distribution of the initial state $|M\rangle^{\otimes n}$. There are two classical registers, CR1 and CR2. The former stores the phase space samples $b(t)$, and the latter the prior measurement record, as in (a). The measurement device M is replaced by a second sampler $\mathsf{\Sigma^{\prime}}$ that takes as input a phase space point $b(t)$ and a Pauli label $a(t)$, and outputs a new phase space point $b(t+1)$ as well as a measurement outcome $s(t)$. The same information that in the standard quantum mechanical description is carried by the quantum state $|\Psi(t)\rangle$ is in the $\Lambda$ polytope description carried by the bit string $b(t)$.
A summary of this sampling process is given in Section II; see in particular Theorem 1 and Algorithm 1. For now, the gist is displayed in Figure 1. A crucial feature of the $\Lambda$ polytope formalism is that the quantum state |$\ket{\Psi(t)}$ of the system (Figure 1a) is replaced by a bit string $b(t)$ of bounded length (Figure 1b). That bit string $b(t)$ is a valid and accurate representation of the quantum system. It should be noted that, in the end, what needs to be reproduced is the quantum mechanical prediction for the joint distribution of measurement outcomes. For it, the statistical distribution of the bit strings $\{b(t),\,\forall t\}$ matters, not individual values $b(t)$. However, this is the same for the quantum mechanical states |$\ket{\Psi(t)}$. They too are conditioned on prior measurement outcomes, hence probabilistic.
@@
The question of interest for the present work is how much classical information the simulation of quantum computations must track, i.e., What is the length of the bit strings $b(t)$? For example, if it turned out that those bit strings were very long, say exponentially long in the number $n$ of magic states, this would provide a convenient explanation for the hardness of classical simulation of universal quantum computation using $\Lambda$ polytopes. If the information storage itself is inefficient, so is the processing. However, this is not what we find. We find that the bit strings $b(t)$ are short. Specifically, they are of length $O(n^{2})$. Thus, simulation of universal quantum computation based on $\Lambda$ polytopes is a small data problem. The presumed hardness of this simulation must come from the computational hardness of the sampling processes involved, not from moving around large amounts of data.
@@
Naively, the length of the bit strings $b(t)$ labeling vertices of $\Lambda_{n}$ is upper-bounded by $\log_{2}(|\mathcal{V}_{n}|)$, with $|\mathcal{V}_{n}|$ the size of the generalized phase space, i.e., the number of vertices of $\Lambda_{n}$. To date, an upper bound and a lower bound are known for this quantity, namely
@@
The lower bound is by Karanjai, Wallman, and Bartlett [10]. The upper bound comes from the upper bound theorem of polytope theory [11] (see Appendix for details).
The gap between the bounds is extremely wide, consistent with both efficient and inefficient storage of the bit strings $b(t)$. From numerical results it seems that the number of phase space points is growing very rapidly with $n$; for $n=1$ the number of phase space points is 8, for $n=2$ it is 22320, and for $n>2$ we don’t know the precise numbers, but the estimate for $n=3$ is already huge.
The following simple insight is crucial for establishing our main result, $|b(t)|=O(n^{2})$: in the QCM model, for any fixed value $n$, all quantum computations start in the same magic state $|M\rangle^{\otimes n}$. Therefore, the question of interest for classical simulation of QCM using $\Lambda$ polytopes is not “What is the size of the phase space $\mathcal{V}_{n}$?”, but rather “What is the size of the region of $\mathcal{V}_{n}$ that can be reached from the initial magic state $|M\rangle^{\otimes n}$?”. This size can be computed, and it turns out to be small. It implies a new upper bound for the length of the bit strings $b(t)$ that is within a factor of four of the matching lower bound [10]. This is the content of our main result, Theorem 2, presented in Section III below.
II Preliminaries
II.1 Quantum computation with magic states
@@
The measurements come from the $n$-qubit Pauli group $\mathcal{P}_{n}$, the group generated by the Pauli operators $X$, $Y$, $Z$ acting on $n$ qubits. Modding out overall phases we have $\mathcal{P}_{n}/\mathcal{Z}(\mathcal{P}_{n})\cong\mathbb{Z}_{2}^{2n}$, and without loss of generality we can fix a phase convention for the Pauli operators to be
@@
$\forall a=(a_{z},a_{x})\in\mathbb{Z}_{2}^{n}\times\mathbb{Z}_{2}^{n}=:E_{n}$ where the inner product $\innerproduct{a_{z}}{a_{x}}$ is computed mod $4$. The projector corresponding to a Pauli measurement $a\in E_{n}$ yielding outcome $s\in\mathbb{Z}_{2}$ is denoted $\Pi_{a}^{s}:=(1+(-1)^{s}T_{a})/2$. The gates of the model are chosen from the Clifford group—the normalizer of the Pauli group in the unitary group up to overall phases, $\mathcal{C}\ell_{n}=\mathcal{N}(\mathcal{P}_{n})/U(1)$. They are defined by the property that they map Pauli operators to Pauli operators under conjugation.
II.2 The $\Lambda$ polytope model
@@
For any fixed number $n\in\mathbb{N}$ of qubits, $\Lambda_{n}$ is a bounded polytope with a finite number of vertices [9]. We denote the vertices of $\Lambda_{n}$ by $\{A_{\alpha}\;|\;\alpha\in\mathcal{V}_{n}\}$ where $\mathcal{V}_{n}$ is an index set. The hidden variable model is defined by the following theorem.
Theorem 1 (Ref. [2]; Theorem 1).
@@
[…]
@@
5:        update the phase space point according to $\alpha\rightarrowtail g\cdot\alpha$
6:     if a Pauli measurement $a\in E_{n}$ is encountered then
7:        sample $(\beta,s)\in\mathcal{V}_{n}\times\mathbb{Z}_{2}$ according to $q_{\alpha,a}$
8:        return  $s\in\mathbb{Z}_{2}$ as the outcome of the measurement
9:        update the phase space point according to $\alpha\rightarrowtail\beta$
III Main result
In this section we present our main result—the amount of classical data that the simulation procedure Algorithm 1 must track is small. First, in section III.1 we give the result for a simplified version of the computational model wherein we only allow Pauli measurements, no Clifford gates, and in particular, we allow only sequences of independent and commuting Pauli measurements. This simplified model is still universal for quantum computation [13, 14]. Then, in Section III.2 we give a more general statement of the main result where we allow computations consisting of any sequence of Clifford gates and Pauli measurements.
III.1 Simplified case
@@
First, in QCM, we usually assume that every computation starts from a fixed magic input state. For example, using the standard magic state circuit gadget [15, Figure 10.25], any Clifford+$T$ circuit with $n$ $T$-gates acting on $m$ qubits can be turned into a QCM circuit (Clifford gates and Pauli measurements only) on $n+m$ qubits acting on a state of the form |$\ket{0}^{\otimes m}\otimes\ket{H}^{\otimes n}$ where |$\ket{H}=(\ket{0}+\exp(i\pi/4)\ket{1})/\sqrt{2}$. In general, it suffices to consider input states of the form |$\ket{0}^{\otimes m}\otimes\ket{M}^{n}$ where |$\ket{M}$ is a fixed single-qubit magic state [16].
@@
We can also do away with the stabilizer part of the input. The Pauli circuit on the $m+n$-qubit state |$\ket{0}^{\otimes m}\otimes\ket{M}^{\otimes n}$ can be simulated by an adaptive sequence of Pauli measurements acting only on the $n$-qubit magic part of the input [17, 14]. The stabilizer part is handled by extra classical side processing.
Finally, it suffices to consider only sequences of commuting Pauli measurements up to length $n$ [14]. This is because when a Pauli measurement is encountered which anticommutes with a measurement that was previously performed, the outcome of the measurement will be uniformly random, and the update of the state after the measurement can be implemented by a Clifford gate. Therefore, the anticommuting measurement can always for replaced by a coin flip to determine the measurement outcome and a Clifford gate to implement the state update which can then be propagated past the future measurements. The longest sequences of independent pair-wise commmuting Pauli measurements on $n$ qubits have length $n$.
@@
5:     update the phase space point according to $\alpha_{t-1}\rightarrowtail\alpha_{t}$
@@
The proof of this theorem relies on the following result from convex geometry (see e.g. Ref. [18, §1.6]).
@@
Proof of Theorem 2. Since the generalized phase space point operators $\{A_{\alpha}\;|\;\alpha\in\mathcal{V}_{n}\}$ are not a basis for $\text{Herm}_{1}(\mathcal{H}_{n})$, they are overcomplete, the distributions $p_{\rho}$ that represent states in the model of Theorem 1, and the distributions $q_{\alpha,a}$ that represent Pauli measurements are not unique. Since $\Lambda_{n}$ lives in $\text{Herm}_{1}(\mathcal{H}_{n})$, a real affine space of dimension $14^{n}-$ 1, by Carathéodory’s theorem there exist choices for the distribution $p_{\rho}$ such that $|\text{supp}(p_{\rho})|\leq 4^{n}$. Similarly, for each $s\in\mathbb{Z}_{2}$, there exist choices for $q_{\alpha,a}(-,s)$ such that $|\text{supp}(q_{\alpha,a}(-,s))|\leq 4^{n}$. To start, we fix a canonical choice of the probability distributions $p_{\rho}$ and $q_{\alpha,a}$ satisfying these properties. With this canonical choice for the distribution representing the input state of the circuit, specifying a sample from this distribution requires no more than $\log_{2}(4^{n})=2n$ bits.
There are $14^{n}-$ 1 nontrivial $n$-qubit Pauli measurements, therefore, specifying each measurement requires no more than $2n$ bits. For the $t^{th}$ measurement $a_{t}$, the distribution $q_{\alpha_{t-1},a_{t}}$ is uniquely specified by the sampling history consisting of states $\alpha_{0},\alpha_{1},\dots,\alpha_{t-1}$, measurements $a_{1},a_{2},\dots,a_{t-1}$, and measurement outcomes $s_{1},s_{2},\dots,s_{t-1}$. Once the distribution is fixed, with the canonical choice above, specifying a sample from this distribution requires no more than $12n+$ 1 bits ($1$ bit for $s_{t}$ and $2n$ bits for $\alpha_{t}$).
Since the length of the measurement sequence is no more than $n$, the number of classical bits required to specify the complete sampling history is no more than
@@
This initial bound can be improved in a number of ways. First, for the purpose of simulation, we don’t need to store $\alpha_{n}$ since there are no more measurements. This immediately removes $2n$ bits. Second, after the $t^{th}$ measurement, the value of any measurement in the span of $a_{1},a_{2},\dots,a_{t}$ is already determined. Therefore, if the $t+1^{th}$ measurement $a_{t+1}$ is to be independent and commute with the previous measurements, it is chosen from $(E_{n}\cap\{a_{1},a_{2},\dots,a_{t}\}^{\perp})/\text{span}(a_{1},a_{2},\dots,a_{t})\cong E_{n-t}$. Specifying a measurement chosen from this set requires only $2(n-t)$ bits, not the full $2n$ bits.
Finally, we can perform another simplification which reduces the number of qubits by $1$ after each measurement. After a measurement of $a\in E_{n}$ giving outcome $s\in\mathbb{Z}_{2}$, the relevant state space is projected to $\Pi_{a}^{s}\Lambda_{n}\Pi_{a}^{s}$. This is contained in a $14^{n-1}-$ 1 dimensional subspace of $\text{Herm}_{1}(\mathcal{H}_{n})$. There exists a Clifford gate $g\in\mathcal{C}\ell_{n}$ such that $g\Pi_{z_{n}}^{0}g^{\dagger}=\Pi_{a}^{s}$, and $\Pi_{a}^{s}\Lambda_{n}\Pi_{a}^{s}=g\Pi_{z_{n}}^{0}\Lambda_{n}\Pi_{z_{n}}^{0}g^{\dagger}=g(\Lambda_{n-1}\otimes\ket{0}\bra{0})g^{\dagger}$. Therefore, after the measurement the Clifford gate $g$ can be propagated out yielding a Pauli measurement circuit on an input state with a stabilizer state tensor factor. Then the above mentioned circuit simplification [17, 14] can be used to remove the stabilizer part of the input. This reduction can be performed after each measurement. This is similar to the idea behind the reduced classical simulation of Theorem 3 of Ref. [17]. With this dimension reduction after each measurement, to specify the sample $(\alpha_{t},s_{t})$ requires no more than $12(n-t)+$ 1 bits. This dimension reduction implicitly implements both of the reductions of the previous paragraph. Note that this explicit dimension reduction is not required to get the memory savings since, according to Ref. [17, Theorem 2], $\Lambda_{n-t}\otimes\ket{0}\bra{0}^{\otimes t}$ is a subpolytope of $\Lambda_{n}$ contained in a $14^{n-t}-$ 1-dimensional affine subspace of $\text{Herm}_{1}(\mathcal{H}_{n})$, and so by Carathéodory’s theorem, an operator of the form $X\otimes\ket{0}\bra{0}^{\otimes t}$ can be decomposed in vertices of this subpolytope with support of size no more than $4^{n-t}$.
With these reductions, the complete measurement history can be specified with no more than
@@
Figure 2: Three descriptions of a quantum computation. (a) A circuit level description. Pauli measurements $a_{1},a_{2},\dots$ are performed on the input state $\rho$ yielding measurement outcomes $s_{1},s_{2},\dots$. (b) The representation of this computation in the model of Theorem 1. The input state is represented by the probability distribution $p_{\rho}$, each measurement $a_{t}$ is represented by a probability distribution $q_{\alpha_{t-1},a_{t}}$. (c) A single run of the simulation algorithm based on the probabilistic representation of the computation. We start by sampling from $p_{\rho}$ to obtain $\alpha_{0}$. For each measurement $a_{i}$ we sample from $q_{\alpha_{t-1},a_{t}}$ to obtain measurement outcome $s_{t}$ and updated state $\alpha_{t}$. (d) An upper bound on the number of classical bits required to store each piece of the sampling history of (c) according to Theorem 2.
A similar idea as the one used in the proof of Theorem 2 of explicitly tracking the measurement history of a quantum circuit has previously been used to define a contextual hidden variable model for the stabilizer subtheory [19].
III.2 General case
Theorem 2 used a simplified version of QCM, where, without loss in computational power, all Clifford gates have been eliminated and the measurement sequence is shrunk to at most $n$ commuting Pauli measurements. In the remainder of this section we demonstrate that the bound on the size of the reachable phase space region does not increase much if we do not make these simplifications, i.e., if we admit arbitrarily long sequences of (potentially non-commuting) Pauli measurements, and Clifford gates between them. Specifically, we have the following additional result.

### v3

@@ I Introduction
This description of quantum computation with magic states grew out of the analysis of Wigner function negativity as a precondition for a quantum computational speedup, a research programme that started with Refs. [3] and [4]. Specifically, in Ref. [4] it was established that negativity in Gross’ Wigner function [5, 6] is required for a quantum speedup, under the condition that the Hilbert space dimension is odd. Analogous results were subsequently established in even dimension, specifically for rebits [7] and qubits [8]. However, in the end it turned out that once sufficiently general (quasi)probability functions are admitted, there is no need for any negativity at all [2, 9]. Universal quantum computation in the magic state model can be described by repeated sampling from a generalized phase space whose points are labeled by the vertices of the $\Lambda$ polytopes. This process essentially resembles a random walk, with the complication that the transition function changes from one time step to the next and can depend on the prior sampling history. A summary of this sampling process is given in Section III; see in particular Theorem 1 and Algorithm 1.
@@
The question of interest for the present work is how much classical information the simulation of quantum computations must track, i.e., What is the length of the bit strings produced in the simulation? For example, if it turned out that those bit strings were very long, say exponentially long in the number $n$ of magic states, this would provide a convenient explanation for the hardness of classical simulation of universal quantum computation using $\Lambda$ polytopes. If the information storage itself is inefficient, so is the processing. However, this is not what we find. We find that the bit strings are short. Specifically, they are of length $O(n^{2})$. Thus, simulation of universal quantum computation based on $\Lambda$ polytopes is a small data problem. The presumed hardness of this simulation must come from the computational hardness of the sampling processes involved, not from moving around large amounts of data. [^1] Elsewhere in the quantum domain there are sampling problems which are algorithmically hard classically but for which samples can be obtained quantum mechanically, for example, boson sampling [26], random circuit sampling [27], and certain instantaneous quantum circuits [28].
@@
Naively, the length of the bit strings labeling the vertices of $\Lambda_{n}$ are upper-bounded by $\log_{2}(|\mathcal{V}_{n}|)$, with $|\mathcal{V}_{n}|$ the size of the generalized phase space, i.e., the number of vertices of $\Lambda_{n}$. To date, an upper bound and a lower bound are known for this quantity, namely
@@
The lower bound is by Karanjai, Wallman, and Bartlett [11]. The upper bound comes from the upper bound theorem of polytope theory [12] (see Appendix A for details).
The gap between the bounds is extremely wide, consistent with both efficient and inefficient storage of the bit strings. Numerical studies suggest that the number of phase space points is growing very rapidly with $n$; for $n=1$ the number of phase space points is 8, for $n=2$ it is 22320, and for $n>2$ we don’t know the precise numbers, but the estimate for $n=3$ is already huge.
The following simple insight is crucial for establishing our main result: in the QCM model, for any fixed value $n$, all quantum computations start in the same magic state $|M\rangle^{\otimes n}$. Therefore, the question of interest for classical simulation of QCM using $\Lambda$ polytopes is not “What is the size of the phase space $\mathcal{V}_{n}$?”, but rather “What is the size of the region of $\mathcal{V}_{n}$ that can be reached from the initial magic state $|M\rangle^{\otimes n}$?”. This size can be computed, and it turns out to be small. It implies a new upper bound for the length of the bit strings that is within a factor of four of the matching lower bound [11]. This is the content of our main result, Theorem 2, presented in Section IV below.
Figure 1: Quantum computation with magic states (a), and its simulation based on $\Lambda$ polytopes (b).
II Quantum state as a string of bits
A crucial feature of the $\Lambda$ polytope formalism is that the quantum state |$\ket{\Psi(t)}$ of the system is replaced by a bit string $b(t)$ of bounded length. That bit string $b(t)$ is a valid and accurate representation of the quantum system.
The role of the bit string $b(t)$ is illustrated in Fig. 1. QCM consists of preparing a quantum register in a magic state |$\ket{M}^{\otimes n}$, followed by a sequence of Pauli measurements; see Fig. 1a. This requires a device Prep to deliver the magic states to the quantum register QR, and a classical register CR to store the previous measurement record s, a classical side computation to identify the label $a(t)$ of the Pauli observable measured in step $t$, and a measurement device M to perform the measurements and to output the corresponding results $s(t)$. The overall structure of the classical simulation is the same, but with the components modified; see Fig. 1b. Prep is replaced by a first sampler $\mathsf{\Sigma}$ that samples from the phase space distribution of the initial state |$\ket{M}^{\otimes n}$. There are two classical registers, CR1 and CR2. The former stores the phase space samples $b(t)$, and the latter the prior measurement record, as in (a). The measurement device M is replaced by a second sampler $\mathsf{\Sigma^{\prime}}$ that takes as input a phase space point $b(t)$ and a Pauli label $a(t)$, and outputs a new phase space point $b(t+1)$ as well as a measurement outcome $s(t)$. The same information that in the standard quantum mechanical description is carried by the quantum state $|\Psi(t)\rangle$ is in the $\Lambda$ polytope description carried by the bit string $b(t)$.
For universal quantum computation in the magic state model, the classical simulation needs to reproduce the quantum mechanical prediction for the joint distribution of outcomes for any sequence of Pauli measurements. The $\Lambda$-polytope method does that, without any approximation. For the predictions, the statistical distribution of the bit strings $\{b(t),\,\forall t\}$ matters, not individual values $b(t)$. This is the same for the quantum mechanical states |$\ket{\Psi(t)}$. They too are conditioned on prior measurement outcomes, hence probabilistic.
III Preliminaries
III.1 Quantum computation with magic states
@@
The measurements come from the $n$-qubit Pauli group $\mathcal{P}_{n}$, the group generated by the Pauli operators $X$, $Y$, $Z$ acting on $n$ qubits. Modding out overall phases we have $\mathcal{P}_{n}/\mathcal{Z}(\mathcal{P}_{n})\cong\mathbb{Z}_{2}^{2n}$, where $\mathcal{Z}(\mathcal{P}_{n})\cong\langle 1,i,-1,-i\rangle$ is the centre of the group $\mathcal{P}_{n}$, and without loss of generality we can fix a phase convention for the Pauli operators to be
@@
$\forall a=(a_{z},a_{x})\in\mathbb{Z}_{2}^{n}\times\mathbb{Z}_{2}^{n}=:E_{n}$ where the inner product $\innerproduct{a_{z}}{a_{x}}$ is computed mod $4$. The projector corresponding to a Pauli measurement $a\in E_{n}$ yielding outcome $s\in\mathbb{Z}_{2}$ is denoted $\Pi_{a}^{s}:=(1+(-1)^{s}T_{a})/2$. The gates of the model are chosen from the Clifford group—the normalizer of the Pauli group in the unitary group up to overall phases, $\mathcal{C}\ell_{n}=\mathcal{N}(\mathcal{P}_{n})/U(1)$. They are defined by the property that they map Pauli operators to Pauli operators under conjugation. For more background on QCM, see Appendix B.
III.2 The $\Lambda$ polytope model
@@
[…]
Theorem 1 (Ref. [2], Theorem 1 & Ref. [9], Theorem 1).
@@
    Any $n$-qubit quantum state $\rho\in\text{Herm}_{1}^{\succeq 0}(\mathcal{H}_{n})$ can be decomposed as
@@
5:        update the phase space point according to $\alpha\leftarrowtail g\cdot\alpha$
6:     end if
7:     if a Pauli measurement $a\in E_{n}$ is encountered then
8:        sample $(\beta,s)\in\mathcal{V}_{n}\times\mathbb{Z}_{2}$ according to $q_{\alpha,a}$
9:        return  $s\in\mathbb{Z}_{2}$ as the outcome of the measurement
10:        update the phase space point according to $\alpha\leftarrowtail\beta$
11:     end if
12:  end while
IV Main result
In this section we present our main result—the amount of classical data that the simulation procedure Algorithm 1 must track is small. First, in section IV.1 we give the result for a simplified version of the computational model wherein we only allow Pauli measurements, no Clifford gates, and in particular, we allow only sequences of independent and commuting Pauli measurements. This simplified model is still universal for quantum computation [15, 16]. Then, in Section IV.2 we give a more general statement of the main result where we allow computations consisting of any sequence of Clifford gates and Pauli measurements.
IV.1 Simplified case
@@
First, in QCM, we usually assume that every computation starts from a fixed magic input state. For example, using the standard magic state circuit gadget [17, Figure 10.25], any Clifford+$T$ circuit with $n$ $T$-gates acting on $m$ qubits can be turned into a QCM circuit (Clifford gates and Pauli measurements only) on $n+m$ qubits acting on a state of the form |$\ket{0}^{\otimes m}\otimes\ket{H}^{\otimes n}$ where |$\ket{H}=(\ket{0}+\exp(i\pi/4)\ket{1})/\sqrt{2}$. In general, it suffices to consider input states of the form |$\ket{0}^{\otimes m}\otimes\ket{M}^{\otimes n}$ where |$\ket{M}$ is a fixed single-qubit magic state [18].
@@
We can also do away with the stabilizer part of the input. The Pauli circuit on the $(m+n)$-qubit state |$\ket{0}^{\otimes m}\otimes\ket{M}^{\otimes n}$ can be simulated by an adaptive sequence of Pauli measurements acting only on the $n$-qubit magic part of the input [19, 16]. The stabilizer part is handled by extra classical side processing.
Finally, it suffices to consider only sequences of commuting Pauli measurements up to length $n$ [16]. This is because when a Pauli measurement is encountered which anticommutes with a measurement that was previously performed, the outcome of the measurement will be uniformly random, and the update of the state after the measurement can be implemented by a Clifford gate. Therefore, the anticommuting measurement can always be replaced by a coin flip to determine the measurement outcome and a Clifford gate to implement the state update which can then be propagated past the future measurements. The longest sequences of independent pair-wise commuting Pauli measurements on $n$ qubits have length $n$.
@@
5:  end for
@@
Note that the precise form of the magic input state used is not important for the proof of this theorem, so long as the state is fixed for each number of qubits. For universal quantum computation, it suffices to consider $n$ copies of a fixed single-qubit magic state such as |$\ket{H}^{\otimes n}$. The proof of this theorem relies on the following result from convex geometry (see e.g. Ref. [20, §1.6]).
@@
Proof of Theorem 2. Since the generalized phase space point operators $\{A_{\alpha}\,|\,\alpha\in\mathcal{V}_{n}\}$ are not a basis for $\text{Herm}_{1}(\mathcal{H}_{n})$, they are overcomplete, the distributions $p_{\rho}$ that represent states and the distributions $q_{\alpha,a}$ that represent Pauli measurements in the model of Theorem 1 are not unique. Since $\Lambda_{n}$ lives in $\text{Herm}_{1}(\mathcal{H}_{n})$, a real affine space of dimension $14^{n}-$ 1, by Carathéodory’s theorem there exist choices for the distribution $p_{\rho}$ such that $|\text{supp}(p_{\rho})|\leq 4^{n}$. Similarly, for each $s\in\mathbb{Z}_{2}$, there exist choices for $q_{\alpha,a}(-,s)$ such that $|\text{supp}(q_{\alpha,a}(-,s))|\leq 4^{n}$. To start, we fix a canonical choice of the probability distributions $p_{\rho}$ and $q_{\alpha,a}$ satisfying these properties. With the distribution $p_{\rho}$ representing the input state of the circuit fixed, we can label the elements of the support by the numbers $1,2,\dots,4^{n}$, and so specifying a sample from this distribution (i.e., specifying an element of the support of $p_{\rho}$), amounts to specifying an integer between $1$ and $4^{n}$, which takes no more than $\log_{2}(4^{n})=2n$ bits.
There are $14^{n}-$ 1 nontrivial $n$-qubit Pauli measurements, therefore, specifying each measurement requires no more than $2n$ bits. For the $t^{th}$ measurement $a_{t}$, the distribution $q_{\alpha_{t-1},a_{t}}$ is uniquely specified by the sampling record consisting of states $\alpha_{0},\alpha_{1},\dots,\alpha_{t-1}$, measurements $a_{1},a_{2},\dots,a_{t-1}$, and measurement outcomes $s_{1},s_{2},\dots,s_{t-1}$. Once this distribution is fixed, with the canonical choice above, specifying a sample from this distribution requires no more than $12n+$ 1 bits ($1$ bit for $s_{t}$, and as before $2n$ bits for $\alpha_{t}$).
Since the length of the measurement sequence is no more than $n$, the number of classical bits required to specify the complete sampling record is no more than
@@
See Figure 2 for a summary of the accounting used to get this bound.
This initial bound can be improved in a number of ways. First, for the purpose of simulation, we don’t need to store $\alpha_{n}$ since there are no more measurements. This immediately removes $2n$ bits. Second, after the $t^{th}$ measurement, the value of any measurement in the span of $a_{1},a_{2},\dots,a_{t}$ is already determined. Therefore, if the $t+1^{th}$ measurement $a_{t+1}$ is to be independent and commute with the previous measurements, it is chosen from $\{a_{1},a_{2},\dots,a_{t}\}^{\perp}/\text{span}(a_{1},a_{2},\dots,a_{t})\cong E_{n-t}$. Specifying a measurement chosen from this set requires only $2(n-t)$ bits, not the full $2n$ bits.
Finally, we can perform another simplification which reduces the number of qubits by $1$ after each measurement. After a measurement of $a\in E_{n}$ giving outcome $s\in\mathbb{Z}_{2}$, the relevant state space is projected to $\Pi_{a}^{s}\Lambda_{n}\Pi_{a}^{s}$. This is contained in a $14^{n-1}-$ 1 dimensional subspace of $\text{Herm}_{1}(\mathcal{H}_{n})$. There exists a Clifford gate $g\in\mathcal{C}\ell_{n}$ such that $g\Pi_{z_{n}}^{0}g^{\dagger}=\Pi_{a}^{s}$, and $\Pi_{a}^{s}\Lambda_{n}\Pi_{a}^{s}=g\Pi_{z_{n}}^{0}\Lambda_{n}\Pi_{z_{n}}^{0}g^{\dagger}=g(\Lambda_{n-1}\otimes\ket{0}\bra{0})g^{\dagger}$. Therefore, after the measurement the Clifford gate $g$ can be propagated out yielding a Pauli measurement circuit on an input state with a stabilizer state tensor factor. Then the above mentioned circuit simplifications [19, 16] can be used to remove the stabilizer part of the input. This reduction can be performed after each measurement. This is similar to the idea behind the reduced classical simulation of Theorem 3 of Ref. [19]. With this dimension reduction after each measurement, to specify the sample $(\alpha_{t},s_{t})$ requires no more than $12(n-t)+$ 1 bits. This dimension reduction implicitly implements both of the reductions of the previous paragraph. Note that this explicit dimension reduction is not required to get the memory savings since, according to Ref. [19, Theorem 2], $\Lambda_{n-t}\otimes\ket{0}\bra{0}^{\otimes t}$ is a subpolytope of $\Lambda_{n}$ contained in a $14^{n-t}-$ 1-dimensional affine subspace of $\text{Herm}_{1}(\mathcal{H}_{n})$, and so by Carathéodory’s theorem, an operator of the form $X\otimes\ket{0}\bra{0}^{\otimes t}$ can be decomposed in vertices of this subpolytope with support of size no more than $4^{n-t}$.
With these reductions, the complete measurement record can be specified with no more than
@@
Figure 2: Three descriptions of a quantum computation. (a) A circuit level description. Pauli measurements $a_{1},a_{2},\dots$ are performed on the input state $\rho$ yielding measurement outcomes $s_{1},s_{2},\dots$. (b) The representation of this computation in the model of Theorem 1. The input state is represented by the probability distribution $p_{\rho}$, each measurement $a_{t}$ is represented by a probability distribution $q_{\alpha_{t-1},a_{t}}$. (c) A single run of the simulation algorithm based on the probabilistic representation of the computation. We start by sampling from $p_{\rho}$ to obtain $\alpha_{0}$. For each measurement $a_{t}$, we sample from $q_{\alpha_{t-1},a_{t}}$ to obtain measurement outcome $s_{t}$ and updated state $\alpha_{t}$. (d) An upper bound on the number of classical bits required to store each piece of the sampling record of (c) according to Theorem 2.
A similar idea as the one used in the proof of Theorem 2 of explicitly tracking the measurement record of a quantum circuit has previously been used to define a contextual hidden variable model for the stabilizer subtheory [21].
IV.2 General case
Theorem 2 used a simplified version of QCM, where, without loss in computational power, all Clifford gates have been eliminated and the measurement sequence is shrunk to at most $n$ commuting Pauli measurements. In the remainder of this section we demonstrate that the bound on the size of the reachable phase space region does not increase much if we do not make these simplifications, i.e., if we admit arbitrarily long sequences of (potentially non-commuting) Pauli measurements, and Clifford gates between them.
Specifically, we have the following additional result.